yagirldezz99:

What is the area of sector BEC when CE = 9 yd?

dude:

Include the image with the sector

yagirldezz99:

its not letting me

dude:

rip `Attach file` feature does not work?

yagirldezz99:

1 attachment
yagirldezz99:

got it

dude:

Alright so 

yagirldezz99:

Find the area of the parallelogram that has a base of 4 m and a height of 5.5 m.

dude:

lol no, I only gave you the equation of how to find it, did you understand it?

yagirldezz99:

yes i answered it , the answer was 27 ft

yagirldezz99:

but this is a different question

dude:

How did you get 27?

dude:

You were supposed to use \(\large \frac{\theta}{360º}\pi r^2\) We know that \(\theta=120º\) and \(r=9\) So we just substitute

dude:

So... \(\large{ \frac{120}{360º}}\small\pi 9^2\\ \frac{1}{3}\pi 81\\ =>27\pi~or~84.823\) Depends on whether is asks in terms of pi

boyyy:

multiply it by 3